# 25mm25mm30mm30mm

### v1

@@ 2. The Sweedler duality for Hom-algebras
3. The Sweedler duality for Hom-module
In this section, we present the structure of Hom-(co)modules and the morphisms of Hom-(co)modules (see Definition 4.1 in [20]). Subsequently, the Sweedler duality of Hom-module is discussed (see [8]).
@@ 3. The Sweedler duality for Hom-module
Let $(G,\mu,\alpha)$ be a Hom-algebra. A right Hom-module of $G$ is a triple $(M,\psi,\gamma)$ where $M$ is a $\mathbb{K}$-linear space, a map $\psi:M\otimes G\rightarrow M$ with $\psi(m\otimes g)=m\cdot g$ and a map $\gamma:M\rightarrow M$ satisfying:
@@
Let $(C,\Delta,\beta)$ be a Hom-coalgebra. A right Hom-comodule of $C$ is a triple $(A,\phi,\epsilon)$ where $A$ is a $\mathbb{K}$-linear space, a map $\phi:A\rightarrow A\otimes C$ with $\psi(a)=\sum_{(a)}a_{(0)}\otimes a_{(1)}$ and a map $\gamma:M\rightarrow M$ satisfying:
@@
The Sweedler duality of the right Hom-module $M$ based on the universal duality $M^{*}$ is defined by
@@
Let $U$ and $V$ be two $\mathbb{K}$-linear space. Denote $\rho:U^{*}\otimes V^{*}\rightarrow(U\otimes V)^{*}$ with corresponding rule $\rho(f\otimes g)(x\otimes y)=f(x)g(y)$, for any $f\in U^{*}$, $g\in V^{*}$, $x\in U$ and $y\in V$. If $\operatorname{dim}U<+\infty$ or $\operatorname{dim}V<+\infty$, $\rho$ is the isomorphism of the linear spaces.
@@
Let $(G,\mu,\alpha)$ be a finite-dimensional Hom-algebra and $(M,\psi,\gamma)$ be a right Hom-module of $G$. Then, $(M^{*},\psi^{*},\gamma^{*})$ is the right Hom-comodule of the Hom-coalgebra $(G^{*},\mu^{*},\alpha^{*})$, where $\psi^{*}:M^{*}\rightarrow M^{*}\otimes G^{*}$ and $\gamma^{*}:M^{*}\rightarrow M^{*}$.
@@
By applying Lemma 2.4, we can get $(G^{*},\mu^{*},\alpha^{*})$ is a Hom-coalgebra. Then we want to show that $\psi^{*}\circ\gamma^{*}=(\gamma^{*}\otimes\alpha^{*})\circ\psi^{*}$. For any $m^{*}\in M$, $m\in M$ and $g\in G$, we have
@@
Then it is sufficient for us to verify that $(\psi^{*}\otimes\alpha^{*})\circ\psi^{*}=(\gamma^{*}\otimes\mu^{*})\circ\psi^{*}$. For any $a^{*}\in M^{*}$ and $x\in M,y,z\in G$,
@@
Now, we will construct the structure of Hom-comodule via the Sweedler duality.
@@
$M^{\circ}$ is the subspace of $M^{*}$.
@@
Firstly, it can be observed that $M^{\circ}$ is closed under $\mathbb{K}$-multiplication. So it is sufficient for us to show that $M^{\circ}$ is closed under addition.
Let $f,g\in M^{\circ}$, there exist finite-codimensional ideals $I$ and $J$ of $G$ such that $f(M\cdot I)=0$ and $g(M\cdot J)=0$. By Theorem 1, we can show that $I\cap J$ is an ideal of $G$ and $G/(I\cap J)$ is finite-dimensional. Then, we have $(f+g)(M\cdot(I\cap J))=0$. And we may conclude that $f+g\in M^{\circ}$, thereby completing the proof. ∎
@@
Let $(M,\psi,\gamma)$ be a right Hom-module of a Hom-algebra $(G,\mu,\alpha)$ where $\alpha$ is an automorphism map of $G$, then its Sweedler duality $(M^{\circ},\psi^{\circ},\gamma^{\circ})$ is a right Hom-comodule of $G^{\circ}$ where $\psi^{\circ}=\psi^{*}|_{M^{\circ}}$, $\gamma^{\circ}=\gamma^{*}|_{M^{\circ}}$ and $\psi^{*}:M^{*}\rightarrow(M\otimes G)^{*}$.
@@
Noticing $f\in M^{\circ}$, we verify that $\psi^{\circ}(f)\in M^{\circ}\otimes G^{\circ}$. There exists a finite-codimensional ideal $J$ of $G$ such that $f(M\cdot J)=0$. Since $\alpha$ is the automorphism map of $G$, $\gamma(M)\subseteq M$ and $JG\subseteq J$, we have
@@
Then we denote $\pi:M\otimes G\rightarrow M\otimes G/(M\cdot J\otimes G+M\otimes J)$ and $\operatorname{Ker}(\pi)=M\cdot J\otimes G+M\otimes J\subseteq\operatorname{Ker}(f\circ\psi).$
By Lemma 2.6, we can induce a unique morphism $\overline{f\psi}:M\otimes G/(M\cdot J\otimes G+M\otimes J)\rightarrow\mathbb{K}$ satisfying $\overline{f\psi}\circ\pi=f\circ\psi$, that is to say
@@
Denote $d_{i}^{*}=\overline{d_{i}}^{*}\circ\pi_{1}$ and $e_{j}^{*}=\overline{e_{j}}^{*}\circ\pi_{2}$. Since $d_{i}^{*}(M\cdot J)=0$ and $\operatorname{Ker}(e_{j}^{*})\supseteq J$, we have $d_{i}^{*}\in M^{\circ}$ and $e_{j}^{*}\in G^{\circ}$. Thus we can show that
@@
since $\gamma(M\cdot J)=\gamma(M)\cdot\alpha(J)\subseteq M\cdot J$. Then we have $\gamma^{\circ}(m)(M\cdot J)=0$, which implies $\gamma^{\circ}(m)\in M^{\circ}$. Thus, $\gamma^{\circ}(M^{\circ})\subseteq M^{\circ}$. Then by Lemma 3.5 and the relationship between the Sweedler duality and the universal duality we can show that $\psi^{\circ}\circ\gamma^{\circ}=(\gamma^{\circ}\otimes\alpha^{\circ})\circ\psi^{\circ}$ and $(\psi^{\circ}\otimes\alpha^{\circ})\circ\psi^{\circ}=(\gamma^{\circ}\otimes\Delta)\circ\psi^{\circ}$. Hence, $M^{\circ}$ is the right Hom-comodule of $G^{\circ}$. ∎
In the subsequent proof, we will study the properties of Hom-comodule $(M^{\circ},\psi^{\circ},\gamma^{\circ})$ proposed in Theorem 4. Based on the above considerations, we continue to derive the morphism of Hom-comodules as follows.
@@
Let $(M,\psi,\gamma)$ and $(N,\psi^{\prime},\gamma^{\prime})$ be two right Hom-modules over Hom-algebra $(G,\mu,\alpha)$ where $\alpha$ is an automorphism map of $G$, $(M^{\circ},\psi^{\circ},\gamma^{\circ})$ and $(N^{\circ},\psi^{\prime\circ},\gamma^{\prime\circ})$ be two corresponding Sweedler dualities over $G^{\circ}$. Let $\sigma:M\rightarrow N$ be a morphism of right Hom-modules. Note that $\sigma^{\circ}:=\sigma^{*}|_{N^{\circ}}:N^{\circ}\rightarrow M^{*}$. Then $\sigma^{\circ}$ is a morphism of Hom-comodules over $G^{\circ}$
@@
since $\sigma(M)\subseteq N$ and $\alpha(J)\subseteq J$. Hence $\sigma^{\circ}(N^{\circ})\subseteq M^{\circ}$. Then we will prove that $(\sigma^{\circ}\otimes\alpha^{\circ})\circ\psi^{\prime\circ}=\psi^{\circ}\circ\sigma^{\circ}$. For any $b^{*}\in N^{\circ}$ and $x\in M,y\in G$, we have
@@
Thus we obtain $(\sigma^{\circ}\otimes\alpha^{\circ})\circ\psi^{\prime\circ}=\psi^{\circ}\circ\sigma^{\circ}$.
Then we vertify $\gamma^{\circ}\circ\sigma^{\circ}=\sigma^{\circ}\circ\gamma^{\prime\circ}$. For all $c^{*}\in N^{\circ}$ and $x\in M$,
$\langle\gamma^{\circ}\circ\sigma^{\circ}(c^{*}),x\rangle$ $=\langle c^{*},\sigma\circ\gamma(x)\rangle$
$=\langle c^{*},\gamma^{\prime}\circ\sigma(x)\rangle$
$=\langle\sigma^{\circ}\circ\gamma^{\prime\circ}(c^{*}),x\rangle.$
Thus it follows that $\gamma^{\circ}\circ\sigma^{\circ}=\sigma^{\circ}\circ\gamma^{\prime\circ}$. Therefore this theorem is proved. ∎

### v2

@@ 2. Sweedler duality for Hom-algebras
Using the Sweedler duality that we constructed, ordinary Hom-associative algebras naturally get an infinite-dimensional correspondence. We present two as examples in the following.
Example 2.12 (Polynomial Hom-algebra).
Let $\mathbb{K}$ be a field and we consider the Hom-algebra $(\mathbb{K}[x],\cdot,\alpha)$ where $\mathbb{K}[x]$ is the set of the polynomials, $\alpha$ is a linear map acting on the indeterminate $x$ with the property $\alpha(x^{m})=\alpha(x)^{m}$ for any $m\in N$ and the bilinear operator $\cdot:\mathbb{K}[x]\rightarrow\mathbb{K}[x]$ with the notation
$x^{m}\cdot x^{n}=\alpha(x^{m})\alpha(x^{n})=\alpha(x)^{\,m+n}\quad(m,n\geq 0).$
A linear functional $\varphi\in A^{*}=\operatorname{Hom}_{\mathbb{K}}(\mathbb{K}[x],\mathbb{K})$ lies in the Sweedler duality
$A^{\circ}\;=\;\bigl{\{}\varphi\in A^{*}\mid\exists\,N\geq 0\leavevmode\nobreak\ |\leavevmode\nobreak\ \varphi\!\bigl{(}(x^{N+1})\bigr{)}=0\bigr{\}}$
iff it annihilates some finite-codimensional ideal $(x^{N+1})$. Hence $(A^{\circ},\Delta,\mu)$ is the Hom-coalgebra induced by the Sweedler duality where the coordinate functionals span the map $\mu$
$\mu(d_{n})(x^{m})=d_{n}(\alpha(x^{m}))=d_{n}(\alpha(x)^{m})=\delta_{n,m},\qquad n,m\geq 0,$
dualising the multiplication $\cdot\colon\mathbb{K}[x]\!\otimes\!\mathbb{K}[x]\to\mathbb{K}[x]$ yields a comultiplication
$\Delta(d_{n})\;=\;\sum_{i=0}^{n}\mu(d_{i})\otimes\mu(d_{\,n-i}),\qquad n\geq 0.$
Example 2.13 (Tensor Hom-algebra $T(V)$ on a finite-dimensional vector space).
Let $V$ be a finite-dimensional vector space over a field $\mathbb{K}$. Its tensor algebra
$T(V)=\mathbb{K}\oplus\bigoplus_{i=1}^{\infty}V^{\otimes i}=\mathbb{K}\;\oplus\;V\;\oplus\;(V^{\otimes 2})\;\oplus\;\cdots$
is a Hom-algebra $(T(V),\cdot,\alpha)$ where the linear map $\alpha:T(V)\rightarrow T(V)$ with the notation $\alpha(V^{\otimes m})=\alpha(V)^{\otimes m}$and the multiplication $\cdot:T(V)\otimes T(V)\rightarrow T(V)$ are given by tensor Hom-concatenation such that for any $a,b\in\mathbb{K}$ and $A\in V^{\otimes p},B\in V^{\otimes q},C\in V^{\otimes r}$,
$A\cdot B=\alpha(A)\otimes\alpha(B),$
$(aA+bB)\cdot C=a(A\cdot C)+b(B\cdot C).$
For any word $w=k\otimes v_{1}\otimes\cdots\otimes v_{n}\,(v_{i}\in V)$ write $|w|=n+1$. Define the evaluation functionals
$d_{w}\colon T(V)\longrightarrow\mathbb{K},\qquad d_{w}(w^{\prime})=\delta_{w,w^{\prime}},$
which clearly annihilate the finite-codimensional ideal $I_{n}=\bigoplus_{m>n}V^{\otimes m}.$ These $d_{w}$ form a basis of the Sweedler duality $(A^{\circ},\Delta,\gamma)$ where the map $\gamma$ with the calculation
$\gamma(d_{w})(aA+bB)=d_{w}(a\alpha(A)+b\alpha(B))=a\delta_{w,p}+b\delta_{w,q},$
dualising the multiplication $\cdot:T(V)\otimes T(V)\rightarrow T(V)$ again,
$\Delta(d_{w})\;=\;\sum_{uv=w}\mu(d_{u})\otimes\mu(d_{v}),$
where the sum runs over every decomposition of $w$ into a prefix $u$ and a suffix $v$.
3. Sweedler duality for Hom-modules
In this section, we will introduce the structure of Hom-(co)modules and the morphisms of Hom-(co)modules (see definition 4.1 in [29]). Based on these structures, the Sweedler duality of Hom-(co)modules will also be treated. In subsequent discourse, we utilize $\cdot$ to represent the algebraic subset Cartesian product, maintaining equivalence to unsigned operations, thereby clarifying operational sequencing in intricate computations, such as $M\cdot J=MJ$.
@@ 3. Sweedler duality for Hom-modules
Let $(G,\mu,\alpha)$ be a Hom-algebra. A right Hom-module of $G$ is a triple $(M,\psi,\gamma)$ where $M$ is a linear space, a map $\psi:M\otimes G\rightarrow M$ with $\psi(m\otimes g)=m\cdot g$ and a map $\gamma:M\rightarrow M$ satisfying:
@@
Let $(C,\Delta,\beta)$ be a Hom-coalgebra. A right Hom-comodule of $C$ is a triple $(A,\phi,\epsilon)$ where $A$ is a linear space, a map $\phi:A\rightarrow A\otimes C$ with $\psi(a)=\sum_{(a)}a_{(0)}\otimes a_{(1)}$ and a map $\epsilon:C\rightarrow C$ satisfying:
@@
The Sweedler duality of the right Hom-module $M$ based on the ordinary duality $M^{*}$ is defined by
@@
Let $U$ and $V$ be two linear spaces. Denote $\rho:U^{*}\otimes V^{*}\rightarrow(U\otimes V)^{*}$ with corresponding rule $\rho(f\otimes g)(x\otimes y)=f(x)g(y)$, for any $f\in U^{*}$, $g\in V^{*}$, $x\in U$ and $y\in V$. If $\operatorname{dim}U<+\infty$ or $\operatorname{dim}V<+\infty$, $\rho$ is an isomorphism of linear spaces.
@@
Let $(G,\mu,\alpha)$ be a finite-dimensional Hom-algebra and $(M,\psi,\gamma)$ be a right Hom-module of $G$. Then, $(M^{*},\psi^{*},\gamma^{*})$ is a right Hom-comodule of a Hom-coalgebra $(G^{*},\mu^{*},\alpha^{*})$, where $\psi^{*}:M^{*}\rightarrow M^{*}\otimes G^{*}$ and $\gamma^{*}:M^{*}\rightarrow M^{*}$.
@@
Applying Lemma 2.4, we can find $(G^{*},\mu^{*},\alpha^{*})$ is a Hom-coalgebra. For any $m^{*}\in M^{*}$, $m\in M$ and $g\in G$, we have
@@
Then it is sufficient for us to check that $(\psi^{*}\otimes\alpha^{*})\circ\psi^{*}=(\gamma^{*}\otimes\mu^{*})\circ\psi^{*}$. For any $a^{*}\in M^{*}$ and $x\in M,y,z\in G$,
@@
Now, we will construct the structure of the Hom-comodule via the Sweedler duality procedure.
@@
Let $G$ be a Hom-algebra and $M$ be a Hom-module. Then $M^{\circ}$ is the subspace of $M^{*}$.
@@
Firstly, it can be observed that $M^{\circ}$ is closed under multiplication. So it is sufficient for us to show that $M^{\circ}$ is closed under addition.
Let $f,g\in M^{\circ}$, there exist finite-codimensional ideals $I$ and $J$ of $G$ such that $f(M\cdot I)=0$ and $g(M\cdot J)=0$. By Theorem 1, we have $I\cap J$ is an ideal of $G$ and $G/(I\cap J)$ is finite-dimensional. Then, we have $(f+g)(M\cdot(I\cap J))=0$. And we may conclude that $f+g\in M^{\circ}$, thereby completing the proof. ∎
@@
Let $(M,\psi,\gamma)$ be a right Hom-module of a Hom-algebra $(G,\mu,\alpha)$ where $\alpha$ is an automorphism map of $G$, then under the Sweedler duality constrcution, $(M^{\circ},\psi^{\circ},\gamma^{\circ})$ is a right Hom-comodule of $G^{\circ}$ where $\psi^{\circ}=\psi^{*}|_{M^{\circ}}$, $\gamma^{\circ}=\gamma^{*}|_{M^{\circ}}$ and $\psi^{*}:M^{*}\rightarrow(M\otimes G)^{*}$.
@@
Noticing that $f\in M^{\circ}$, we aim to verify that $\psi^{\circ}(f)\in M^{\circ}\otimes G^{\circ}$. There exists a finite-codimensional ideal $J$ of $G$ such that $f(M\cdot J)=0$. Since $\alpha$ is the automorphism map of $G$, $\gamma(M)\subseteq M$ and $JG\subseteq J$, we have
@@
Denote $\pi:M\otimes G\rightarrow M\otimes G/(M\cdot J\otimes G+M\otimes J)$ and $\operatorname{Ker}(\pi)=M\cdot J\otimes G+M\otimes J\subseteq\operatorname{Ker}(f\circ\psi).$ Using Lemma 2.6, we can induce a unique morphism $\overline{f\psi}:M\otimes G/(M\cdot J\otimes G+M\otimes J)\rightarrow\mathbb{K}$ satisfying $\overline{f\psi}\circ\pi=f\circ\psi$, that is to say,
@@
Denote $d_{i}^{*}=\overline{d_{i}}^{*}\circ\pi_{1}$ and $e_{j}^{*}=\overline{e_{j}}^{*}\circ\pi_{2}$. Since $d_{i}^{*}(M\cdot J)=0$ and $\operatorname{Ker}(e_{j}^{*})\supseteq J$, we have $d_{i}^{*}\in M^{\circ}$ and $e_{j}^{*}\in G^{\circ}$. Thus, we have
@@
since $\gamma(M\cdot J)=\gamma(M)\cdot\alpha(J)\subseteq M\cdot J$. Then we have $\gamma^{\circ}(m)(M\cdot J)=0$, which implies $\gamma^{\circ}(m)\in M^{\circ}$. Thus, $\gamma^{\circ}(M^{\circ})\subseteq M^{\circ}$. Combining Theorem 1, Theorem 3, and Lemma 3.5 together, we deduce that $\psi^{\circ}\circ\gamma^{\circ}=(\gamma^{\circ}\otimes\alpha^{\circ})\circ\psi^{\circ}$ and $(\psi^{\circ}\otimes\alpha^{\circ})\circ\psi^{\circ}=(\gamma^{\circ}\otimes\Delta)\circ\psi^{\circ}$. Hence, $M^{\circ}$ is a right Hom-comodule of $G^{\circ}$. ∎
Remark 3.6.
In the proof of Theorem 4 we must insist that the twisting map $\alpha:G\to G$ is bijective. Surjectivity gives $\alpha(G)=G$ and, in particular, $\alpha(J)=J$ for every ideal $J\subseteq G$. Hence
$M\!\cdot\!J\otimes G\;=\;M\!\cdot\!J\otimes\alpha(G),$
which implies
$\psi^{\circ}(f)\bigl{(}M\!\cdot\!J\otimes G+M\otimes J\bigr{)}=0.$
This step is essential for establishing that $\psi^{\circ}(f)$ lies in $M^{\circ}\otimes G^{\circ}$.
In the following, we will study the properties of the Hom-comodule $(M^{\circ},\psi^{\circ},\gamma^{\circ})$ proposed in Theorem 4. Based on the above considerations, we continue to derive the morphism of Hom-comodules as follows.
@@
Let $(M,\psi,\gamma)$ and $(N,\psi^{\prime},\gamma^{\prime})$ be two right Hom-modules over Hom-algebra $(G,\mu,\alpha)$ where $\alpha$ is an automorphism map of $G$, $(M^{\circ},\psi^{\circ},\gamma^{\circ})$ and $(N^{\circ},\psi^{\prime\circ},\gamma^{\prime\circ})$ be two corresponding objects under the Sweedler duality over $G^{\circ}$. Let $\sigma:M\rightarrow N$ be a morphism of right Hom-modules. Note that $\sigma^{\circ}:=\sigma^{*}|_{N^{\circ}}:N^{\circ}\rightarrow M^{*}$. Then $\sigma^{\circ}$ is a morphism of Hom-comodules over $G^{\circ}$.
@@
since $\sigma(M)\subseteq N$ and $\alpha(J)\subseteq J$. Hence $\sigma^{\circ}(N^{\circ})\subseteq M^{\circ}$.
For any $b^{*}\in N^{\circ}$ and $x\in M,y\in G$, we have
@@
Thus we obtain $(\sigma^{\circ}\otimes\alpha^{\circ})\circ\psi^{\prime\circ}=\psi^{\circ}\circ\sigma^{\circ}$. And then for all $c^{*}\in N^{\circ}$ and $x\in M$,
$\langle\gamma^{\circ}\circ\sigma^{\circ}(c^{*}),x\rangle=\langle c^{*},\sigma\circ\gamma(x)\rangle=\langle c^{*},\gamma^{\prime}\circ\sigma(x)\rangle=\langle\sigma^{\circ}\circ\gamma^{\prime\circ}(c^{*}),x\rangle.$
Thus it follows that $\gamma^{\circ}\circ\sigma^{\circ}=\sigma^{\circ}\circ\gamma^{\prime\circ}$. ∎